If * (x, y) = 30(y + y ^ 2) represents the population density of a planar region on Earth, where randy are measured in miles, find the number of people in the region bounded by the curves x = y ^ 2, x = 2y - y ^ 2


1
Expert's answer
2021-11-01T17:37:59-0400

When do the curves "x=y^2" and "x=2y-y^2" intersect?


"y^2=2y-y^2"

"2y(y-1)=0"

"y_1=0, y_2=1"

Express the number of people in the region as iterated integral


"\displaystyle\int_{0}^{1}\displaystyle\int_{y^2}^{2y-y^2}30(y+y^2)dxdy"

"=\displaystyle\int_{0}^{1}30(y+y^2)[x]\begin{matrix}
 2y-y^2 \\
 y^2
\end{matrix}dy"


"=\displaystyle\int_{0}^{1}30(y+y^2)(2y-2y^2)dy"

"=\displaystyle\int_{0}^{1}60(y^2-y^4)dy"

"=[20y^3-12y^5]\begin{matrix}
 1\\
 0
\end{matrix}"

"=20-12=8"

The number of people in the region bounded by the curves is "8."